What is the sum of the solutions to the equation 9 x^2 - 6x = 8?
Answer by mananth(16946)

9 x^2 - 6x = 8?
9x^2-6x-8=0
9x^2-12x+6x-8=0
3x(3x-4)+2(3x-4)=0
(3x+2)(3x-4)=0
..
3x=-2
x= -2/3
or
3x-4=0
3x=4
x=4/3
..
sum of solutions = -2/3 +4/3
=2/3